OpenStudy (anonymous):

indicate whether the following are acidic, basic, or neutral solution A. [H2O+]=6.0x10^-12 B. [H3O+]=1.4x10^-9 C. [OH-]=5.0x10^-12 D. {OH-]=3.5x10^-10

OpenStudy (anonymous):

Check out their pH. from these formula's pH=-log [H+] pOH=-log[OH-] pH=14-pOH of 1st and 2nd use the formula pH=-log[H+] to calculate pH. for 3rd and 4th use pOH=-log[OH-] to calculate pOH, then from it calculate pH. i.e pH=14-pOH pH > 7 = Basic pH=7=Neutral pH<7=Acidic

OpenStudy (anonymous):

I feel like i am doing something wrong putting the formula in the calculator for A I keep getting .0004

OpenStudy (anonymous):

-log(6.0x10^12)

OpenStudy (anonymous):

A. pH=-log(6.0X10^-12)=11.2 i.e Base B. pH=-log(1.4X10^-9)=8.8 i.e Base C. pOH=-log(5.0X10^-12)=11.3 pH=14-pOH = 14-11.3= 2.7 = Acidic D. pOH=-log(3.5X-10)=9.45 pH=14-pOH=14-9.45=4.55=Acidic

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

welcome